OpenStudy (anonymous):

You deposit $20 at the end of every three months into an annuity with 4% interest compounded quarterly. How much of the annuity's value is from interest after 25 years? Do not round until the final answer. Then round to the nearest dollar as needed.

OpenStudy (here_to_help15):

Use the Future Value of an Ordinary Annuity equation: FVoa = PMT [((1 + i)^n - 1) / i], where i = the quarterly rate 0.04/4 = 0.01, n = 25 yrs * 4 qtrs per year = 100, PMT is the quarterly deposit of $20... FVoa = 20 [(1.01^100) - 1 / 0.01] = 20 [(2.70481 - 1) / 0.01] = 20 [170.48138] = 3,409.63....now subtract your payments to get the amount of interest:... (20 * 100) = 2,000..interest is 1409.63, round to 1410

OpenStudy (anonymous):

Thank you for the help! Now I understand.

OpenStudy (here_to_help15):

You don't do you?

OpenStudy (here_to_help15):

Do you need a lower explanation?

OpenStudy (anonymous):

Could you please give me one? That would help.

OpenStudy (here_to_help15):

yes